# Pre-trained Encoder Inference: Revealing Upstream Encoders In Downstream Machine Learning Services

### v1

@@ VIII Case Study: PEI-assisted Adversarial Attacks Against LLaVA
Figure 10: Two examples of adversarial attacks against LLaVA with adversarial iamges synthesized based on the hidden image encoder revealed by the PEI attack. The synthesized adversarial images contain visually benign mosaics, but can induce LLaVA to generate predefined false health/medical information.
Overview of the attack. We consider an adversarial attack that aims to use visually benign adversarial images to induce LLaVA to generate harmful information. For instance, as shown in Figure VII, although the adversarial images only contain visually benign mosaics, LLaVA can read out misleading health/medical information from them. These adversarial images can be used to stealthily spread harmful information, as their harmfulness is difficult for humans to detect. To launch such an adversarial attack with the assistance of the PEI attack, the adversary will first use the PEI attack to reveal the hidden image encoder (i.e., CLIP-ViT-L/14-336px) used by LLaVA-1.5-13B. Then, the adversary will synthesize adversarial examples directly based on the hidden CLIP image encoder. We then discuss the two stages of the attack in detail.
Results of the PEI attack against LLaVA. We consider a PEI candidates set consists of seven image encoders, which are: CLIP ViT-B/16 [2], CLIP ViT-B/32, CLIP ViT-L/14, CLIP ViT-L/14-336px, OpenCLIP ViT-B/32 [62], OpenCLIP ViT-H/14, and OpenCLIP ViT-L/14. For the synthesis of PEI attack images, we adopt the same objective images and hyperparameters described in Section VI-A. For the PEI score calculation, to evaluate the behavior similarity in Eq. (2), we propose to leverage the question-answering capability of LLaVA to directly score the similarity between objective image and PEI attack images (in the range $[0,1]$). See Appendix B-A for the detailed design of the PEI score calculation.
The PEI attack results are presented in Table VIII, in which the correct hidden encoder, i.e., CLIP Vit-L/14-336px, is the only one that enjoys a $z$-score above the preset threshold $1.7$. That means our PEI attack successfully reveals the image encoder hidden in LLaVA-1.5-13B.
Results of the PEI-assisted adversarial attack against LLaVA. After the image encoder (i.e., CLIP ViT-L/14-336px) used by LLaVA-1.5-13B is revealed by the PEI attack, we then synthesize adversarial images that can induce LLaVA generating harmful contents solely based on the revealed image encoder. Specifically, as shown in Figure 2, to launch the attack, we first construct a targeted image that contains predefined harmful plain text content. Then, since the revealed hidden CLIP image encoder is an open-source model, we can synthesize an adversarial image by minimizing its embedding difference from that of the targeted image in a white-box manner. This will result in an adversarial image that contains benign mosaics but LLaVA can read out the predefined harmful content from it. Please refer to Appendix B-B for more details about the adversarial image synthesis.
We present two attack examples in Figure 2, in which one can find that LLaVA reads out false medical/health information from PEI-assisted adversarial images. This verifies the effectiveness of the PEI-assisted adversarial attack and further demonstrates the usefulness and hazards of the PEI attack.
IX Potential Defenses
IX-A Transforming PEI Attack Samples
A possible direction of defending the PEI attack is to perform data transformation on PEI attack samples before feeding them to the protected downstream service to sanitize their harmfulness. While many existing filtering-based methods [65, 66] could be used to remove malicious information in PEI attack samples, they may also reduce model performance on clean data. However, a series of works also focus on making adversarial examples robust to data transformation [67, 68, 69, 70]. Nevertheless, here we focus on analyzing two simple plug-and-play and downstream task-agnostic attack samples sanitizing methods for image and text encoders, respectively.
A plug-and-play preprocessing defense for image encoders. As PEI attack samples can somewhat be seen as adversarial examples [71], one may naturally seek to leverage defenses originally proposed for adversarial attacks to tackle the new PEI attack. As a preliminary investigation, here we adopt a task-agnostic and plug-and-play method named JPEG defense [72, 73] to protect downstream services from the PEI attack. Concretely, for each query image, the service supplier will first process it with the canonical JPEG compressing algorithm before feeding it into the real downstream service $g$. Since JPEG compression is found to be effective in destroying malicious information within adversarial images [72, 73], it may also be able to mitigate hazards of PEI attack images. A case study is conducted on two image classification services, “RN34 (HF) + CIFAR-10” and “RN50 (MS) + CIFAR-10”.
The results of JPEG defense against the vanilla PEI attack are presented in Figure 11, in which we plot $z$-scores of different candidates with and without the JPEG defense (denoted as “Origin” and “JPEG-Def” respectively). From the results, we find that the JPEG defense effectively defeats the PEI attack: in all cases, under the JPEG defense, the PEI $z$-scores of the correct hidden encoder are suppressed to low levels and no candidate goes beyond the preset threshold $1.7$. However, we also discovered a simple method to bypass the JPEG defense for PEI attack: we just rescale each PEI attack image from $64\times 64$ to $128\times 128$ before sending them to the target service. Results are also presented in Figure 11 (denote as “JPEG-Def v.s. Resize-128”), which show that this rescaling approach enables the PEI attack to succeed again in all analyzed cases. We deduce the mechanism behind this is that resizing an attack image to a larger scale can prevent local features that contain malicious information from being compressed by JPEG defense.
(a) RN34 (HF) + CIFAR-10.
(b) RN50 (MS) + CIFAR-10.
Figure 11: PEI $z$-scores of candidates on image classification services. (1) “Origin”: original PEI attack results. (2) “JPEG-Def”: results of PEI attack against JPEG defense. (3) “JPEG-Def v.s. Resize-128”: results of PEI attack strengthened by image resizing (to shape $128\times 128$) against JPEG defense.
A Plug-and-play preprocessing defense for text encoders. As a preliminary investigation, here we analyze a simple character-flipping defense. Specifically, for each input text, the service supplier will randomly select and flip a part of its characters to other (lowercase) alphabets before feeding it to the downstream service $g$. We conduct case studies on SST-5 with encoders BERT (base), BERT (large), and CLIP ViT-L/14, with a flipping proportion of $20\%$. In these three settings, the PEI attack originally succeeds in revealing the correct hidden encoders (see Table II). Results are presented as Figure 12, from which we find that while the random-flipping strategy effectively defeats the PEI attack in most settings, it could not always succeed (e.g., in the case “BERT (base) + SST-5”). In addition, while a larger flipping proportion may improve the defense, it may also compromise the qualities of benign inputs.
IX-B Detecting PEI Attack Samples
Another direction of potential defenses is to leverage Out-of-distribution (OOD) detection methods to identify and drop malicious queries toward protected ML services. For image data, a series of detection methods [74, 75, 76, 77, 78] against adversarial images or backdoor images could be applied to defend against the PEI attack. For text data, it is found that perplexity filtering [79] can effectively identify gibberish but malicious texts from human-readable texts, while gibberish text is indeed used in PEI against text encoders.
However, PEI attack samples can also be made stealth to bypass potential detection defenses. For example, for image data, one may be able to exploit detection bypassing methods originally designed for adversarial images [80, 81, 82] to PEI attack images. For text data, by leveraging the strong natural language generation ability of LLMs, the adversary may further synthesize PEI attack texts that enjoy strong human-readability [69, 25, 83] and could thus be made to be robust to gibberish text detection-based defenses.
(a) BERT (base).
(b) BERT (large).
(c) CLIP ViT-L/14.
Figure 12: PEI $z$-scores of candidates on SST-5 text classification services. (1) “Origin”: original PEI attack results. (2) “Flipping-Def (20%)”: results of PEI attack against character flipping defense (with a proportion of $20\%$).
IX-C Architecture Resistance
More reliable defenses may be re-designing the pipeline of building downstream ML services to make them natively robust to the PEI attack. For example, since the current version of the PEI attack assumes that the targeted downstream service is built upon a single encoder, one may leverage multiple upstream encoders to build a single service to bypass the PEI attack. The idea of using multiple pre-trained encoders to build downstream models has already been adopted to improve the downstream model performance (e.g., SDXL [84]). Our results suggest that apart from improving performance, there is also a great need to use multiple pre-trained encoders for the sake of downstream model security.
X Conclusions
This paper unveils a novel privacy attack named pre-trained encoder inference (PEI) attack, which aims to reveal pre-trained encoders that are hidden in downstream ML services. We design a general framework to perform the PEI attack in a black-box manner and further instantiate it for image encoders and text encoders. Extensive experiments on three downstream tasks, i.e., image classification, text classification, and text-to-image generation, empirically verify the effectiveness of our new PEI attack. We further conduct a case study on a recent and advanced multimodal model, i.e., LLaVA, to show the usefulness of the the PEI attack in assisting other ML attacks, such as adversarial attacks, against LLaVA. Results demonstrate the great need to design general defenses against the PEI attack. Additionally, we discuss several potential defenses against the PEI attack.
Appendix A Additional Experiment Details in Section VI
A-A Details of Datasets
A-A1 Downstream Datasets for Image Classification
@@
Four image datasets are used as the downstream data. They are: CIFAR-10 [51] that consists of $50,000$ training images and $10,000$ test images from $10$ classes; STL-10 [53] that consists of $5,000$ training images and $8,000$ test images from $10$ classes; SVHN [52] that consists of $73,257$ training images and $26,032$ test images from $10$ classes; and Food-101 [54] that consists of $75,750$ training images and $25,250$ validation images (used as test data) from $101$ classes.
A-A2 Downstream Datasets for Text Classification
Four text datasets are used as the downstream data. They are: SST-5 [55] consists of $8,544$ training sentences and $1,101$ validation sentences (used as test data) from $5$ classes; Yelp [56] consists of $650,000$ training sentences and $50,000$ test sentences from $5$ classes; AG-News [56] consists of $120,000$ training sentences and $7,600$ test sentences from $4$ classes; and TREC [57] consists of $5,452$ training sentences and $500$ test sentences from $6$ classes.
A-B Objective Samples in the PEI Attack
A-B1 Objective Images
For the PEI attack against image encoders, we randomly sampled $10$ images from the PASCAL VOC 2012 dataset [61] as the objective image samples. They are presented in Figure VII.
Figure 13: The $10$ objective images used in the PEI attack against image encoders.
A-B2 Objective Texts
For the PEI attack against text encoders, we manually constructed $20$ sentences following the template {Adjective} + {Noun} as the objective text samples. They are listed in Table V.
TABLE V: The $20$ objective texts used in the PEI attack against text encoders.
1 black bridge red sunflower 11
2 dark battery dirty battery 12
3 orange cherry black mobilephone 13
4 yellow rabbit heavy carcamera 14
5 white network green aircraft 15
6 hot fish muddy album 16
7 dirty softdrink purple network 17
8 green album hot house 18
9 beautiful oven yellow bridge 19
10 purple dragon ice opera 20
TABLE VI: Test accuracy (%) of downstream image classification services built upon different pre-trained encoders.
Downstream Dataset Upstream Pre-trained Encoder
RN34 (HF) RN50 (HF) RN34 (MS) RN50 (MS) MobileV3 CLIP ViT-L/14
CIFAR-10 91.24 91.88 90.44 91.05 91.34 98.11
SVHN 62.93 67.22 60.99 69.67 71.23 82.58
STL-10 97.16 97.86 96.74 97.92 96.44 99.83
Food-101 64.40 68.66 63.26 68.67 71.20 94.66
TABLE VII: Test accuracy (%) of downstream text classification services built upon different pre-trained encoders.
Downstream Dataset Downstream Pre-trained Encoder
BERT (base) BERT (large) T5 (small) T5 (base) RoBERTa (base) CLIP ViT-L/14
SST-5 40.51 43.14 40.69 43.42 43.87 39.15
TREC 92.20 85.00 93.40 94.40 94.00 91.80
Yelp 51.20 46.52 55.28 59.86 60.00 55.08
AG-News 89.46 79.71 91.39 93.07 93.30 93.07
TABLE VIII: Pricing data of common real-world EaaSs.
Type Supplier Model Name Price Link
Image Encoder Vertex AI (by Google) multimodalembeddings \$0.0001 / Image Input https://cloud.google.com/vertex-ai/generative-ai/pricing Accessed Date: 2024-04
Azure AI (by Microsoft) Image Embeddings \$0.10 / 1,000 Transactions https://azure.microsoft.com/en-us/pricing/details/cognitive-services/computer-vision/ Accessed Date: 2024-04
Text Encoder OpenAI text-embedding-3-small \$0.02 / 1M Tokens https://openai.com/pricing Accessed Date: 2024-04
OpenAI text-embedding-3-large \$0.13 / 1M Tokens https://openai.com/pricing Accessed Date: 2024-04
OpenAI ada v2 \$0.10 / 1M Tokens https://openai.com/pricing Accessed Date: 2024-04
Vertex AI (by Google) multimodalembeddings \$0.0002 / 1K Characters Input https://cloud.google.com/vertex-ai/generative-ai/pricing Accessed Date: 2024-04
Azure AI (by Microsoft) Text Embeddings \$0.014 / 1,000 Transactions https://azure.microsoft.com/en-us/pricing/details/cognitive-services/computer-vision/ Accessed Date: 2024-04
@@ A-B2 Objective Texts
TABLE IX: Download links of all pre-trained encoders and downstream models adopted in our experiments.
@@
Image Encoder ResNet-34 (HF) https://huggingface.co/timm/resnet34.a1_in1k
@@
Text Encoder BERT (base) https://huggingface.co/google-bert/bert-base-uncased
BERT (large) https://huggingface.co/google-bert/bert-large-uncased
T5 (small) https://huggingface.co/google-t5/t5-small
T5 (base) https://huggingface.co/google-t5/t5-base
RoBERTa (base) https://huggingface.co/FacebookAI/roberta-base
@@
CLIP Model CLIP ViT-B/16 https://huggingface.co/openai/clip-vit-base-patch16
@@
Text-to-image Model Stable Diffusion-v1.2 https://huggingface.co/CompVis/stable-diffusion-v1-2
Stable Diffusion-v1.4 https://huggingface.co/CompVis/stable-diffusion-v1-4
Stable Diffusion-base-v2.1 https://huggingface.co/stabilityai/stable-diffusion-2-1-base
Stable Diffusion-v2.1 https://huggingface.co/stabilityai/stable-diffusion-2-1
@@
A-C Downstream utilities of classification tasks
We collect and present the classification accuracy of downstream image/text classification services built on different downstream data and pre-trained encoders. The results of downstream image classification services are reported in Table A-B2, while those of downstream text classification services are reported in Table A-B2.
Appendix B Additional Experiment Details in Section VIII
B-A PEI Score Calculation
@@
$\ell_{sim}(g(x^{(atk)}_{i,j,k}),g(x^{(obj)}_{j}))$
$:=\frac{1}{2}\left(\ell_{ask}(x^{(atk)}_{i,j,k},x^{(obj)}_{j})+\ell_{ask}(x^{(obj)}_{j},x^{(atk)}_{i,j,k})\right).$
B-B PEI-assisted Adversarial Example Synthesis
To synthesize adversarial examples based on the PEI-revealed image encoder (i.e., CLIP ViT-L/14-336px), we propose to minimize their embedding difference with that of a targeted image containing preset harmful text content. Specifically, we aim to minimize the squared loss defined by embeddings of the adversarial image and targeted image. Since the hidden CLIP model is an open-source model, one can further leverage first-order gradient descent to optimize the objective squared loss. We use sign gradient to minimize the objective loss for $2,000$ iterations, in which the learning rate is fixed to $0.01$. Besides, to further improve the success rate of obtaining valid PEI-assisted adversarial images, for each targeted image, we will first synthesize $16$ candidate adversarial images, and then choose a single image that enjoys the best attack performance among them as the eventual output.
Appendix C Pre-trained Models Download Links
All pre-trained encoders and downstream models in our experiments are available on the Hugging Face Model Repository. We list the download links to these models in Table A-B2.

### v2

@@ 5.2 Case study: PEI-assisted adversarial attack
Figure 3: Two examples of adversarial attacks against LLaVA with adversarial images synthesized based on the hidden vision encoder revealed by the PEI attack. These adversarial images contain visually benign mosaics, but can induce LLaVA to generate predefined false health/medical information.
6 Conclusions
We unveiled a new pre-trained encoder inference (PEI) attack that can threaten both deployed pre-trained encoders and downstream ML services without modifying the service-building pipeline. The goal of the PEI attack is to reveal encoders hidden in targeted downstream services. Such encoder information can then be used to threaten original targeted services. We conducted experiments on vision encoders and found that the PEI attack can effectively: (1) reveal encoders hidden in different downstream services, and (2) facilitate other ML attacks against targeted services. Our results call for the need to design general defenses against the new PEI attack to protect downstream ML services.
Appendix A Related works
Privacy attacks against pre-trained encoders. For data privacy attacks, many efforts have been made to design MIA against encoders. When the pre-training strategy is known, liu2021encodermi [43] successfully performs membership inference attacks (MIA) against targeted encoder based on membership scores calculated with the pre-training loss function and shadow training techniques. he2022semi [30] extends MIA against semi-supervisedly learned encoders. zhu2024unified [74] studies a more challenging setting where the adversary has no knowledge about pre-training algorithms and proposes to calculate the embedding similarity between global and local features as the membership score for each sample. Besides, for model privacy attacks, studies mainly focus on stealing the powerful encoding functionality of encoders. By simply querying the targeted encoder and collecting return embeddings, liu2022stolenencoder [45] succeeds in retraining an encoder that reproduces the functionality of the target with a reasonable query budget. sha2023can [58] adopted a contrastive learning-based method to further improve the encoder stealing performance. However, these data/model privacy attacks can only threaten pre-trained encoders on the upstream side, while our PEI attack can further threaten downstream ML services.
Encoder attacks against downstream models. To increase the vulnerability of downstream models, adversaries usually try to attack upstream encoders by manipulating their pre-training stage with poisoned or backdoored training data. A series of works have studied poisoning/backdooring vision encoders when the downstream task is image classification. carlini2022poisoning [6] and jia2022badencoder [35] showed that by injecting trigger patches to the training data, classifiers built upon the backdoored encoder can effectively misclassify inputs with backdoor triggers to a target label. carlini2022poisoning [6] further found that only backdooring a significantly modest amount of training data (around $0.01\%$) is enough to perform strong attacks. Other advances include liu2022poisonedencoder [42], zhang2022corruptencoder [72] and tao2024distribution [60]. More recent works have started to exploit encoder backdoor attacks to increase the privacy vulnerability of downstream models built upon backdoored encoders, for example, making downstream training data more vulnerable to membership inference attacks [67] or data reconstruction attacks [21]. However, these attacks usually require detailed information about the downstream tasks, such as the number of classes in downstream classification tasks or prior knowledge of potential targeted training data, which may limit their practicality in the real world. As a comparison, our PEI attacks can be performed in a downstream task-agnostic manner without modifying the pre-training stage of upstream encoders.
Appendix B Omitted details of Section 4
This section collects additional experimental details omitted from Section 4.
B.1 Additional experimental setup
@@
Downstream datasets. Three image datasets are used as downstream data: CIFAR-10 [37] that consists of $50,000$ training images and $10,000$ test images from $10$ classes; SVHN [47] that consists of $73,257$ training images and $26,032$ test images from $10$ classes; and Food-101 [2] that consists of $75,750$ training images and $25,250$ validation images (used as test data) from $101$ classes.
Pre-trained encoders. Six vision encoders are adopted as candidates in the PEI attack, which are: ResNet-34 (HF) [29], ResNet-50 (HF), MobileNetV3 [32], ResNet-34 (MS), ResNet-50 (MS), and CLIP ViT-L/14 [54]. All these encoders can be downloaded from the Hugging Face Model Repository. Download links are collected and presented in Table 4.
@@
Table 4: Download links of pre-trained vision encoders adopted in our experiments.
@@
Vision Encoder ResNet-34 (HF) https://huggingface.co/timm/resnet34.a1_in1k
@@
Building downstream classifiers. For each pair of encoder and dataset, we fix the encoder and train a downstream classifier based on embeddings of downstream training data obtained from the encoder. For each service, the classifier is an MLP consisting of three fully-connected layers, where the dimension of each hidden layer is $512$. We use Adam to train the classifier for $10,000$ iterations. The batch size is set as $512$. The learning rate is set as $0.001$ and decayed by $0.1$ every $4,000$ iterations.
B.2 PEI objective images
For the PEI attack against vision encoders, we randomly sampled $10$ images from the PASCAL VOC 2012 dataset [20] as the objective image samples. They are presented in Figure 4.3.
Figure 4: The $10$ objective images used in the PEI attack against vision encoders.
B.3 Downstream utilities of image classification
We collect and present the classification accuracy of downstream image classification services built on different downstream data and pre-trained encoders. The results of downstream image classification services are reported in Table 5.
Table 5: Test accuracy (%) of downstream image classification services built upon different pre-trained encoders.
Downstream Dataset Upstream Pre-trained Encoder
RN34 (HF) RN50 (HF) RN34 (MS) RN50 (MS) MobileNetV3 CLIP ViT-L/14
CIFAR-10 91.24 91.88 90.44 91.05 91.34 98.11
SVHN 62.93 67.22 60.99 69.67 71.23 82.58
Food-101 64.40 68.66 63.26 68.67 71.20 94.66
B.4 Ablation studies of Section 4.2
This section collects ablation studies omitted from Section 4.2.
Effect of downstream classifier initialization. The PEI attack in Table 1 is conducted against single downstream classifiers. Yet, it remains unknown how it would be affected by the random initialization of these classifiers. To investigate this, for each dataset-encoder pair, we retrain $5$ downstream models, perform the PEI attack against all of them, and report rates of attack success, false-negative, and false-positive in Figure 5.
From the figure, we have two observations. Firstly, PEI attacks never made false-positive inferences in all analyzed cases. As explained in the previous section, this capability is useful for PEI adversaries to quickly find victims from a large number of targeted services. Secondly, in most of the cases (except two cases “RN34 (MS) + SVHN”, and “MobileNetV3 + SVHN”), the attack achieved consistent results, i.e., either attack success or false-positive, with a possibility of at least $80\%$, in $16$ out of $18$ analyzed cases. This suggests that the PEI attack is generally insensitive to the initialization of downstream classifiers in image classification services.
Figure 5: Rates of different PEI attack results on CIFAR-10, SVHN, and Food-101 datasets. Rates of attack success, false-negative, and false-positive are colored differently.
Effect of downstream classifier architectures. So far, the architecture of downstream classifiers used in image classification services is fixed to a $3$-layer MLP with a width of $512$. We now analyze how the architecture of this classifier would affect the PEI attack performance. The analysis is conducted on two cases: “RN34 (HF) + CIFAR-10” and “RN50 (MS) + CIFAR-10”. For each case, we build classification services with the targeted encoder and four different downstream classifiers: MLP-1, MLP-2, MLP-3, and MLP-4. Here, “MLP-$x$” denotes an MLP architecture of $x$ layers, where dimensions of all hidden layers (if it has) are fixed to $512$. The training of these downstream classifiers follows that described in Section B.1. After that, we attack each service with PEI attack samples originally used in Section 4.2. PEI $z$-scores are reported in Figure 6. From the figure, we find that the downstream classifier architecture has little effect on attacks against targeted services: the PEI $z$-scores of each candidate are almost the same across different downstream classifier architectures, and the PEI attack always revealed the correct hidden encoder.
(a) RN34 (HF) + CIFAR-10.
(b) RN50 (MS) + CIFAR-10.
Figure 6: PEI $z$-scores of candidates on classification services with downstream classifiers of $4$ different architectures. $z$-scores of different services are in different colors.
Effect of the zeroth-order gradient estimation on the PEI attack image synthesis. According to Eq. (5), the gradient estimation accuracy depends on the random sampling number $S$. We thus vary $S$ in the set $\{25,50,75,100,120\}$ (in Section 4, $S=100$) to see how the attack performance would change. This study is conducted on “RN34 (HF) + CIFAR-10” and “RN50 (MS) + CIFAR-10”. PEI $z$-scores of different encoders under different $S$ are plotted in Figure 7. For the first case (Figure 6(a)), the sampling number $S$ has little effect on PEI $z$-scores, while the $z$-score of the correct hidden encoder always stays in high level (around $2.0$). For the second case (Figure 6(b)), as $S$ increases, the PEI $z$-score of the correct encoder significantly increases, while those of other candidates remain at low levels. These imply that when the PEI attack is weak, it can benefit from a large sampling number $S$. However, since a large $S$ would increase the query budget, the adversary needs to carefully trade-off between the attack utility and efficiency.
(a) RN34 (HF) + CIFAR-10.
(b) RN50 (MS) + CIFAR-10.
Figure 7: PEI $z$-scores of candidates on image classification services with different gradient estimation sampling number $S$ (see Algorithm 2) varies in $\{25,50,75,100,125\}$.
B.5 Additional experimental details of Section 4.3
This section collects omitted experimental details of the PEI-assisted model stealing in Section 4.3.
Building stolen models. We train every stolen model via Adam for $10,000$ iterations, where the batch size is $64$ and the learning rate is initialized as $0.001$ and decayed by $0.1$ every $4,000$ iterations. The surrogate data is constructed from $50$ thousand images randomly selected from the validation set of ImageNet [14], which results in a query budget of $50$ thousand.
Appendix C Omitted details of Section 5
This section collects additional experimental details omitted from Section 5.
C.1 Pre-trained models download links
The targeted downstream multimodal generative model is LLaVA-1.5-13B [40, 41], which is built upon a finetuned Vicuna-1.5-13B language model [73] and the origional CLIP ViT-L/14-336px vision encoder [54]. We adopt seven vision encoders to form the PEI candidates set, which are: CLIP ViT-B/16 [54], CLIP ViT-B/32, CLIP ViT-L/14, CLIP ViT-L/14-336px, OpenCLIP ViT-B/32 [12], OpenCLIP ViT-H/14, and OpenCLIP ViT-L/14. All these models can be downloaded from the HUgging Face Model Repository. Download links are collected and presented in Table 6.
Table 6: Download links of pre-trained models adopted in Section 5.
Model Type Name Link
@@ C.1 Pre-trained models download links
Vision Encoder CLIP ViT-B/16 https://huggingface.co/openai/clip-vit-base-patch16
@@
C.2 PEI score calculation
@@
$\ell_{sim}(g(x^{(atk)}_{i,j,k}),g(x^{(obj)}_{j})):=\frac{1}{2}\left(\ell_{ask}(x^{(atk)}_{i,j,k},x^{(obj)}_{j})+\ell_{ask}(x^{(obj)}_{j},x^{(atk)}_{i,j,k})\right).$
C.3 PEI-assisted adversarial example synthesis
To synthesize adversarial examples based on the PEI-revealed vision encoder (i.e., CLIP ViT-L/14-336px), we propose to minimize their embedding difference with that of a targeted image containing preset harmful text content. Specifically, we aim to minimize the squared loss defined by embeddings of the adversarial image and targeted image. Since the hidden CLIP model is an open-source model, one can further leverage first-order gradient descent to optimize the objective squared loss. We use sign gradient to minimize the objective loss for $2,000$ iterations, in which the learning rate is fixed to $0.01$. Besides, to further improve the success rate of obtaining valid PEI-assisted adversarial images, for each targeted image, we will first synthesize $16$ candidate adversarial images, and then choose a single image that enjoys the best attack performance among them as the eventual output.